OpenStudy (anonymous):

Given that \(\Large a+b=3\) and \(\Large a^2+b^2=7\), find the value of \(\Large a^{10}+2a^9b+3a^8b^2+4a^7b^3+5a^6b^4\\\Large +6a^5b^5+5a^4b^6+4a^3b^7+3a^2b^8+2ab^9+b^{10}\)

OpenStudy (vishweshshrimali5):

Very interesting question..

OpenStudy (vishweshshrimali5):

First of all using a+b and a^2+b^2 you can find out ab

OpenStudy (vishweshshrimali5):

(a+b)^2 = (a^2+b^2) + 2ab

OpenStudy (vishweshshrimali5):

Then : (a-b)^2 = a^2 + b^2 - 2ab Using a^2+b^2 and ab, find out (a-b)

OpenStudy (vishweshshrimali5):

Then using a-b and a+b values we can find out a and b

OpenStudy (vishweshshrimali5):

Plug in the values and solve...

OpenStudy (vishweshshrimali5):

Well this was a primary school answer.. lets search for an olympiad type answer

OpenStudy (anonymous):

ab=1

OpenStudy (vishweshshrimali5):

Great

OpenStudy (akashdeepdeb):

hehe, I started thinking of binomial theorem.

OpenStudy (vishweshshrimali5):

:D @AkashdeepDeb I too but when I saw the coefficients I knew it would not be the best choice for now

OpenStudy (vishweshshrimali5):

Well we will search for a better answer later on first lets find out the answer.

OpenStudy (vishweshshrimali5):

@JungHyunRan , you have calculated ab. Now try finding out (a-b)

OpenStudy (akashdeepdeb):

Use \[(a-b)^2 = a^2 + b^2 - 2ab\]

OpenStudy (anonymous):

\(\Large a-b=\sqrt{5}\) right?

OpenStudy (akashdeepdeb):

\[(a-b) = \sqrt{a^2 + b^2 - 2ab}\] Yes.

OpenStudy (vishweshshrimali5):

I would have preferred the use of \(\pm\), in such a good question

OpenStudy (anonymous):

ah, yup :) \(\Large a-b=\pm\sqrt{5}\)

OpenStudy (vishweshshrimali5):

\[(a-b) = \pm \sqrt{5}\]

OpenStudy (akashdeepdeb):

Then there'll be multiple values for a and b and thus creating more values for the final answer right?

OpenStudy (vishweshshrimali5):

No then we will use the fact that ab is positive and thus a and b must have the same sign

OpenStudy (vishweshshrimali5):

Generally in such cases, the answers are a and b interchanged

OpenStudy (vishweshshrimali5):

The same would be here :)

OpenStudy (vishweshshrimali5):

Which when you notice the expression would find out that it will not affect whether a and b are interchanged :)

OpenStudy (akashdeepdeb):

Excellent! Thanks @vishweshshrimali5 !

OpenStudy (vishweshshrimali5):

So you can take \(\pm\) or just forget about that :) My pleasure @AkashdeepDeb :)

OpenStudy (vishweshshrimali5):

Well lets find out a and b @JungHyunRan . We have a+b and a-b

OpenStudy (vishweshshrimali5):

Great and b ?

OpenStudy (anonymous):

\(\Large \begin{cases} a=\frac{3+\sqrt{5}}{2} \\ b=\frac{3-\sqrt{5}}{2}\end{cases}\) OR \(\Large \begin{cases}a=\frac{3-\sqrt{5}}{2}\\ b=\frac{3+\sqrt{5}}{2}\end{cases}\)

OpenStudy (vishweshshrimali5):

Great now select any one set because the answer would remain same whichever set you select

OpenStudy (anonymous):

yes

OpenStudy (vishweshshrimali5):

So which one do you select ? 1st one or 2nd ?

OpenStudy (anonymous):

I select 1st

OpenStudy (vishweshshrimali5):

Very good

OpenStudy (vishweshshrimali5):

Now lets see what the expression was

OpenStudy (vishweshshrimali5):

\[\Large a^{10}+2a^9b+3a^8b^2+4a^7b^3+5a^6b^4\\\Large +6a^5b^5+5a^4b^6+4a^3b^7+3a^2b^8+2ab^9+b^{10}\] Now first of all lets simplify it... use the fact that ab = 1

OpenStudy (vishweshshrimali5):

Thus, \(\large{2a^9b = 2a^8 * ab = 2a^8}\) and so on

OpenStudy (vishweshshrimali5):

So your expression would become ?

OpenStudy (anonymous):

\(\Large (a^{10}+b^{10})+ab(2a^8+3a^7b+4a^6b^2+5a^5b^3+6a^4b^4\\\Large +5a^3b^5+4a^2b^6+3ab^7+2b^8)\)

OpenStudy (vishweshshrimali5):

Good then use the fact that ab = 1...

OpenStudy (vishweshshrimali5):

Your expression would become: \[\Large a^{10}+2a^8+3a^6+4a^4+5a^2\\\Large +6+5b^2+4b^4+3b^6+2b^8+b^{10}\]

OpenStudy (vishweshshrimali5):

Looks much better. Doesn't it ?

OpenStudy (anonymous):

yeah :)

OpenStudy (vishweshshrimali5):

\[\large{=(a^{10} + b^{10}) + 2(a^8 + b^8) + 3(a^6 + b^6) + 3(a^4 + b^4) + 5(a^2+b^2) + 6}\]

OpenStudy (vishweshshrimali5):

We also know (a^2+b^2) value so lets plug in that too. What will we get ?

OpenStudy (anonymous):

So, how to find value of \(\Large a^{10}+b^{10}\) ?

OpenStudy (vishweshshrimali5):

Lets start in increasing order.

OpenStudy (vishweshshrimali5):

First lets evaluate \(a^2 + b^2\) but we already know that it is 7

OpenStudy (vishweshshrimali5):

Then lets find out a^4 + b^4

OpenStudy (anonymous):

\(\Large a^4+b^4=(a^2+b^2)^2-2a^2b^2=7^2-2=47\)

OpenStudy (vishweshshrimali5):

Great !!

OpenStudy (vishweshshrimali5):

Then \(\large{a^6+b^6}\) ?

OpenStudy (vishweshshrimali5):

Try using (a^4+b^4) and (a^2+b^2)

OpenStudy (vishweshshrimali5):

Or better use (a^3+b^3)^2

OpenStudy (vishweshshrimali5):

First find out (a^3+b^3): \[\large{a^3+b^3 = (a+b)(a^2+b^2-ab) = 3(7-1) = 18}\]

OpenStudy (anonymous):

\(\Large a^6+b^6=(a^3+b^3)^2-2a^3b^3=18^2-2=322\)

OpenStudy (vishweshshrimali5):

Very good

OpenStudy (vishweshshrimali5):

Now we have : (a^8 + b^8)

OpenStudy (vishweshshrimali5):

Use (a^4+b^4)

OpenStudy (anonymous):

\(\Large =(a^4+b^4)^2-2a^4b^4=47^2-2=2207\)

OpenStudy (vishweshshrimali5):

Great

OpenStudy (vishweshshrimali5):

Now only (a^10 + b^10)

OpenStudy (vishweshshrimali5):

\[(a^5 + b^5) = (a^2 + b^2)(a^3+b^3) - a^2b^3 - b^2a^3 = (a^2+b^2)(a^3+b^3) - a^2 b^2(a+b)\]

OpenStudy (vishweshshrimali5):

Now using this find out (a^5+b^5)

OpenStudy (anonymous):

\(\Large a^5+b^5=7\times 18-3=123\)

OpenStudy (vishweshshrimali5):

Great

OpenStudy (anonymous):

Thus, \(\Large a^{10}+b^{10}=(a^5+b^5)^2-2a^5b^5=123^2-2=15127\)

OpenStudy (vishweshshrimali5):

Fantastic

OpenStudy (vishweshshrimali5):

Now lets simplify final answer

OpenStudy (vishweshshrimali5):

What will we get ?

OpenStudy (anonymous):

20736

OpenStudy (vishweshshrimali5):

Great (I have not checked it using calculator so I am trusting you on it) :)

OpenStudy (vishweshshrimali5):

Fantastic work @JungHyunRan . I enjoyed helping you :)

OpenStudy (anonymous):

Thank you :D

OpenStudy (vishweshshrimali5):

Your welcome :) Good day